one angle of a triangle is 30 degrees larger than the smallest angle, and the third angle is four times as large as the smallest angle. find the angle of the smallest angle
Answer by nerdybill(5404)

one angle of a triangle is 30 degrees larger than the smallest angle, and the third angle is four times as large as the smallest angle. find the angle of the smallest angle
.
First, we know that for any triangle, the sum of the interior angles is 180.
.
Let x = smallest angle (angle two)
then from: "one angle of a triangle is 30 degrees larger than the smallest angle"
x+30 = angle one
and from: "the third angle is four times as large as the smallest angle"
4x = angle three
.
x + x+30 + 4x = 180
30 + 6x = 180
6x = 150
x = 25 degrees (smallest angle)